OpenStudy (anonymous):

Solve for the variable in this equation. 2x/5+3/4=9/10

OpenStudy (anonymous):

x = 9/8

hero (hero):

\[\frac{2x}{5} + \frac{3}{4} = \frac{9}{10}\]

hero (hero):

@Ahmed.53 giving answers only is a violation of the code of conduct. I will report it to the mods.

OpenStudy (anonymous):

And i dont even think 9/8 is the answer can you help me @Hero

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

i blocked him. Will you help me Hero?

hero (hero):

Okay

OpenStudy (anonymous):

Thanks:) If i seem a little slow, its because i havent learned this yet considering i missed a month of school so i need extra tutoring you know?

hero (hero):

Yes, okay

OpenStudy (anonymous):

Thanks

hero (hero):

I will set up vyew and send you a link

OpenStudy (anonymous):

Okay thanks:)

myininaya (myininaya):

@Ahmed.53 Hero is right! Please do not be abusive to others, do not use bad language, and do not give only answers. Please at least provide an explanation. Please review the CoC: http://openstudy.com/terms-and-conditions . Thanks.